### v1

Abstract
This paper considers the need for generalizable bias mitigation techniques in machine learning due to the growing concerns of fairness and discrimination in data-driven decision-making procedures across a range of industries. While many existing methods for mitigating bias in machine learning have succeeded in specific cases, they often lack generalizability and cannot be easily applied to different data types or models. Additionally, the trade-off between accuracy and fairness remains a fundamental tension in the field. To address these issues, we propose a bias mitigation method based on multi-task learning, utilizing the concept of Monte-Carlo dropout and Pareto optimality from multi-objective optimization. This method optimizes accuracy and fairness while improving the model’s explainability without using sensitive information. We test this method on three datasets from different domains and show how it can deliver the most desired trade-off between model fairness and performance. This allows for tuning in specific domains where one metric may be more important than another. With the framework we introduce in this paper, we aim to enhance the fairness-performance trade-off and offer a solution to bias mitigation methods’ generalizability issues in machine learning.
Keywords Bias  $\cdot$ Fairness  $\cdot$ Monte-Carlo dropout  $\cdot$ Pareto Optimality
1 Introduction
The term bias was first introduced in the machine learning domain by Tom Mitchell in his 1980 paper titled “The need for biases in learning generalizations” Mitchell (1980). The concept of bias refers to giving importance to particular features to improve generalization. This general idea of bias in machine learning is positive and necessary for models to perform, eliminating the risk of hyper-focusing on specific samples over others. On the contrary, bias can also be negative in machine learning. Negative bias can be defined as an inaccurate assumption made by a machine learning algorithm that is systematically or historically prejudiced against certain groups of people Zanna et al. (2022). Decisions made by these biased algorithms could cause adverse effects on particular social groups, for example, those defined by sex, race, age, marital status, handicaps, etc., when used to make autonomous decisions in life-changing cases such as health, hiring, education, criminal sentencing, etc.
Negative bias can be introduced into the machine pipeline in two main ways, through the data or the algorithm itself Blanzeisky and Cunningham (2021). Bias due to data, also known as a negative legacy Cunningham and Delany (2021); Kamishima et al. (2012), can be caused by an imbalance in the representation of different population categories in the training data, meaning if certain demographics are lacking from the sample data, models trained on this data often do not generalize when applied to new data that contains those missing demographics Gorrostieta et al. (2019). Another cause of negative legacy is labeling errors. This can occur because data fed into models, especially supervised or semi-supervised models which are widely used in various jurisdictions, are manually labeled by humans who are inherently biased. These models can ultimately reflect people’s impressions and sustain or further magnify bias from the labeled data Vallor (2017). Bias caused by the algorithm itself is regarded as underestimation, and it occurs when the classifier under-predicts an uncommon outcome for the minority group. This can happen when the classifier under-fits the data, causing it to focus on strong signals and missing more subtle phenomena Cunningham and Delany (2021).
For over 10 years, research on bias and fairness in machine learning has been extensive and numerous works have explored the development of analytical techniques that are aware of social fairness and discrimination Kamishima et al. (2012); Friedler et al. (2019). These studies have shown that simply eliminating the sensitive features from the model’s workings is insufficient to avoid an unfavorable determination process. This is due to the indirect influence of sensitive information on other features vital to the algorithm’s decision-making process. This highlights the need and emergence of bias mitigation techniques.
Despite their seeming success in specific cases, there is a recurring trend of bias mitigation methods lacking generalizability. With applications of machine learning in data-driven decision-making procedures in a rapidly growing number of industries and the fairness concerns that arise from them, there is an increasing need for generalizable and more easily transferable bias mitigation techniques. This is because the machine learning algorithms developed and deployed on large scales are usually distinct and often proprietary and have different concerns with bias and limitations when choosing methods to mitigate it. For instance, searches for distinctively Black-sounding names are more likely to trigger ads for arrest records Sweeney (2013). In contrast, image searches for professions such as CEO produce fewer images of women Obermeyer et al. (2019), and natural language processing algorithms encode language in gendered ways Caliskan et al. (2017), among other examples. These are all highly different types of data and models that often cannot be analyzed or manipulated in the same manner. Unfortunately, most existing solutions for mitigating bias and unfairness are one-off applications for a specific problem or use case Pagano et al. (2022). This calls for a need to develop more generalizable bias mitigation techniques that can be implemented with various model types.
Another setback with developing methods to mitigate bias that researchers face is the trade-off between accuracy and fairness, which is prevalent in existing literature Dutta et al. (2020). This challenge proves the need for methods that reconcile this fundamental tension in the bias and fairness aspect of machine learning. As a solution to this issue, several previous works have used the concept of Pareto optimality from multi-objective optimization to seek the fairness-accuracy Pareto front of a neural network classifier Wei and Niethammer (2022); Liu and Vicente (2022); Wang et al. (2021a); Liang et al. (2022); Little et al. (2022); Kamani et al. (2021). Before the introduction of this concept to fairness literature within the past 2 years, existing algorithmic fairness methods have typically performed the so-called linear scalarization scheme, which has limitations in recovering Pareto optimal solutions Wei and Niethammer (2022).
In this work, we propose a bias mitigation method based on multi-task learning (MTL), which has been used in previous research to optimize both accuracy and fairness Oneto et al. (2019), along with Monte Carlo dropout to utilize model uncertainty to improve fairness without a massive sacrifice in either computational complexity or performance. Our method is an extension of our previous work Zanna et al. (2022), a hybrid in-processing and post-processing technique that does not require access to a machine learning model’s objective function and prediction methodology. This method enables us to optimize accuracy and fairness while improving the explainability of the model without explicitly using any of the sensitive information in the functional form of our final model. Researchers have mentioned that analyzing bias and unfairness collaborates with model explainability, so explainability is intrinsic to the study Pagano et al. (2022). We also utilize the concept of Pareto optimality from multi-objective optimization to arrive at models that deliver the most desired trade-off between fairness and performance. This allows the results to be tuned according to the needs in any domain where one metric may be more important than another. We analyze and test our method on two publicly available (ADULT & MIMIC-III) and 1 non-public (SNAPSHOT) dataset from three varying domains: in-hospital mortality, finance, and stress prediction, to demonstrate how transferable the method is between different fields. One such dataset is one of the most well-used data for bias studies (ADULT dataset).
2 Related Works
2.1 Definition of Fairness
The concept of fairness has been around in philosophy and psychology and, recently, in machine learning. Current literature has shown a consensus that fairness has to be studied as the endpoint of implementing algorithms Pleiss et al. (2017). Enforcing fairness when building machine learning models requires that fairness is first defined. Philosophy and psychology have tried to define the concept of fairness long before computer science, and the fact that there is no universally agreed-upon definition of fairness further complicates the problem Saleiro et al. (2018). Generally, fairness can be defined as the absence of prejudice or favoritism towards an individual or a group based on their intrinsic or acquired traits in decision-making Saxena et al. (2019). Even though the necessity of fairness has been demonstrated through numerous perspectives, it can be challenging to achieve in practice in machine learning and society as a whole Mehrabi et al. (2021). In this section, we explore different classes of bias mitigation techniques used to attempt to achieve different definitions of fairness on data from different domains and how effective they are for these cases.
2.2 Bias Mitigation Techniques
When it comes to methods developed to counter the issue of negative bias, there are several techniques Calders and Verwer (2010); Chouldechova (2017); Feldman et al. (2015); Hardt et al. (2016); Kamiran and Calders (2012); Zafar et al. (2017); Zhang et al. (2018) that are generally categorized into 3 groups: pre-processing, in-processing, and post-processing methods. Most of these existing bias mitigation methods try to avoid unethical interference of sensitive (protected) attributes in the functional form of models, while others try to include participants from disadvantaged groups Mehrabi et al. (2021). In addition, some works try to satisfy one or more of the fairness notions in their methods, such as disparate impact (DI) ratios, differences in false negative and true positive scores, etc., to satisfy notions of treatment disparity and impact disparity by allowing the protected attributes during the training phase, but avoiding them at prediction Mehrabi et al. (2021). This section will discuss some of these techniques from recent literature, their weaknesses, and what our proposed method provides to bridge the gaps.
2.2.1 Pre-processing Techniques
Perhaps the easiest bias mitigation methods to develop and implement are pre-processing techniques. These involve eliminating the sensitive attributes, changing the class labels, and reweighing or resampling data to remove imbalances without relabeling instances Kamiran and Calders (2012). Pre-processing techniques are used if the algorithm is allowed to modify the training data. They are generally easier to implement, and there is no need to access sensitive labels at test time, which is not allowed in some jurisdictions d’Alessandro et al. (2017); Yan et al. (2020). A good example of this class of techniques is by the authors of Bryant et al. (2019), who implement a pre-processing method to remove sensitive data (e.g., race, gender, etc.) from the model for a financial banking system. They aimed to ensure customer data protection, anonymization, and fairness without affecting the machine learning model. To achieve this, they generate synthetic data from the representation of the original data in order to preserve privacy while maintaining statistical similarity to the original data using Variational Auto-Encoders (VAE). Their method performed well on some datasets they tested. Another work that utilized the method of data synthesizing is Jang et al. (2021), where the authors generated a new dataset that does not contain disparity of distribution, quality, or noise to ensure that all classes are treated equally by building the model on purely generated data that is balanced (to ensure fairness), and then transferring to real data for fine-tuning (to ensure performance). They also utilized the VAE-GAN to approximate the data distribution and generate data with desired sensitive attributes and labels. Georgopoulos et al. (2021) also utilized the idea of synthesizing data in computer vision. A bias mitigation method was proposed to increase demographic diversity in a facial image dataset using Generative Adversarial Networks (GANs) to create additional images by transferring demographic attributes to each image in the unbalanced dataset. These methods generally improve fairness, but they are very domain-oriented, and the same technique cannot be implemented on various data types. There is also some domain-related limitation where for example with health-related data, generation of synthetic data can be tricky. Chen et al. (2021) states that the roles of synthetic data in AI systems in healthcare remain unclear. Specifically, when generating synthetic medical data for rare or new disease conditions, there may not be sufficient samples to establish clinical reference standards. Some works utilize the reweighing pre-processing technique for bias mitigation Paviglianiti and Pasero (2020); Vega-Gonzalo and Christidis (2022); Krasanakis et al. (2018). This method has shown to work reasonably well in improving fairness, but it usually struggles when the dataset does not contain sufficient samples for a specific demographic group Yang et al. (2022).
Our proposed method which is a hybrid between in-processing and post-processing does not incur the domain-related limitations that pre-processing techniques typically suffer from.
2.2.2 In-processing Techniques
In-processing techniques try to modify and change learning algorithms to remove discrimination during the model training process Gorrostieta et al. (2019) - either by incorporating changes into the objective function or imposing a constraint Liu and Vicente (2022). If it is allowed to change the learning procedure for a machine learning model and the sensitive attributes information is accessible in the training samples, in-processing techniques can be used during the training of a model Yan et al. (2020); Gorrostieta et al. (2019); Oneto et al. (2019). In recent years, in-processing techniques have become more prevalent Pagano et al. (2022) due to them generally resulting in better performance on accuracy and fairness. For example, Radovanović et al. (2020) presents a bias mitigation technique targeting group bias. In this work, they modify the logistic regression algorithm to add fairness constraints into the model learning process, forcing the predictive model not to create disparate impact and allow equal opportunity for every sub-population. The authors also reiterated the advantage of in-processing techniques over pre-processing since they do not require altering the data, which hurts ethical norms.
Another in-processing method is introduced in Ashokan and Haas (2021) where a mitigation strategy in which the algorithm learns the difference between predicted and observed ratings in subgroups. The algorithm identifies which is biased and corrects the prediction. Fairness here improved in most use cases, but performance highly varied. Du et al. (2021) attempts to make Deep Neural Networks (DNNs) fairer by targeting the bias only in the output layer. The authors manipulate the structure of the model and regularize its training, resulting in training samples with different sensitive attributes being neutralized. This method reduced the model’s dependence on the sensitive attribute and subsequently improved fairness. This method performed well in terms of fairness but came with a high performance loss.
Despite these aforementioned methods seemingly delivering all they need in terms of fairness, in-processing techniques are also task-specific since they involve manipulating the model itself, which makes them non-transferable, and sometimes modifying the model may not be possible Oneto et al. (2019). In addition to this, these methods often come at a high performance cost. With our proposed method, we tackle the high performance loss by utilizing the Pareto optimality concept that allows models to be tuned to a particular domain where one metric, either fairness or performance, may be more important than the other. Our method also does not require access to a machine learning model’s objective function and prediction methodology, eliminating the issue of in-processing techniques having to be task-specific or modifying the model.
2.2.3 Post-processing Techniques
These methods aim to edit posteriors in ways that satisfy the fairness constraint, and no changes are made to the model Oneto et al. (2019). Post-processing techniques can be used in a scenario where a machine learning algorithm can only treat the learned model as a black box without any ability to modify the training data or learning algorithm. The advantages of this class of techniques are that they can be applied to any classifier since they eliminate the need for modifications. They also have relatively good performance, especially regarding fairness measures.
As a recent form of post-processing for bias mitigation in models, several prior works Zhang et al. (2018); Li et al. (2022); Grari et al. (2019); Adel et al. (2019) have utilized adversarial learning to mitigate bias. Adversarial learning is a method that aims to trick machine learning models by providing deceptive input Huang et al. (2011). It includes generating and detecting adversarial examples, which are inputs specially created to deceive classifiers. Zhang et al. (2018) presents a framework for mitigating bias by including a variable for the interest group and simultaneously learning a predictor and an adversary. The authors claimed that in theory, the method could enforce the chosen metric across multiple definitions of fairness, regardless of the complexity of the predictor’s model. As a result, they were able to train models that were significantly less biased but could not get these models to converge in the end. As for Li et al. (2022), their work focused on studying the fairness of outlier detection. They propose an outlier detection method that utilizes adversarial networks to learn the optimal representation of the original data while hiding the sensitive attribute in the data. Their method was able to improve group fairness and outlier detection performance. In Grari et al. (2019), Grari et al. developed an approach of adversarial gradient tree boosting where the objective is to predict the output with gradient tree boosting while minimizing the ability of an adversarial neural network to predict the sensitive attribute. This approach achieved a similar level of fairness to other proposed algorithms. And lastly in Adel et al. (2019), the authors made slight adjustments to their architecture, given any existing differentiable classifier by adding a new hidden layer, to enable concurrent adversarial optimization for fairness and accuracy.
The general drawback to post-processing methods is that they do not allow for flexibility with the accuracy-fairness trade-off Oneto et al. (2019); Yan et al. (2020); Gorrostieta et al. (2019), which we aim to provide with our proposed method. Additionally, with the current bias mitigation methods, there exists minimal support for black-box models compared to the abundance of information for white-box models. There is also a need for transparency and explainability in machine learning models and mitigation methods that are transferable across various model structures and data types. With our work, we aim to address some of these concerns.
3 Methods
3.1 Fairness Terminology
In this subsection, we introduce and define some of the concepts and terminologies generally used in algorithmic fairness research that are relevant to this paper.
  • Protected (sensitive) label: An attribute that partitions a population into groups whose outcomes should have parity (such as race, gender, income, etc.).
  • Privileged class: A protected label value indicating a group that is at an advantage. Generally represented by the value 1.
  • Unprivileged class: A protected label value indicating a group that is at a disadvantage. Generally represented by the value 0.
  • Disparate impact ratio (DIR): This is a fairness metric which compares the rate of positive outcomes in the unprivileged class, by the rate of positive outcomes in the privileged class, as shown in equation 1. It is the measure of how different outcomes are for different groups, based on the results of a model Feldman et al. (2015). A value of less than 1 indicates less favorable outcomes for the unprivileged group. In this paper, we assume an acceptable lower bound of 0.8, and an upper bound of 1.2, with 1 being the ideal score.
    DIR =$\text{DIR =}\frac{Pr(Y=1|D=unprivileged)}{Pr(Y=1|D=privileged)}$ (1)
    where Y represents the target label and D the class of the protected demographic label.
  • Equalized odds: This fairness metric enforces that the model correctly identify the positive outcome at equal rates across groups, and misclasify the positive outcome at equal rates across groups (creating the same proportion of True Positives and False Positives across groups). In this work, we represent this metric using 2 values; Difference in False Negative scores, and Difference in False Positive scores.
3.2 Multi-task Learning Based Bias Mitigation Method (Proposed Method)
Our proposed methodology is based on the premise of epistemic uncertainty in Bayesian uncertainty estimation. Uncertainty estimation refers to estimating the confidence level of a model output, and epistemic uncertainty refers to uncertainty in the structure and parameters of a model caused by a lack of knowledge, which has been proven to correlate with model weights Li et al. (2021); Wang et al. (2021b). According to the Bayesian approach, the posterior distribution of model weights can theoretically be obtained and used to make inferences by fitting a defined model to the observed data (training set). In this case, the posterior variance represents the total uncertainty of the prediction Wang et al. (2021b).
We hypothesize that when the model is most uncertain about the protected label, the weights of that model will lack knowledge at certain regions of the feature space related to the protected label. This means that using these weights for target label prediction will minimize the influence of that protected label on the final prediction. Based on this premise, our proposed method will minimize any bias introduced to the model regarding the protected label.
Figure 1 represents a visual diagram of our proposed method, which we described in our initial work (Zanna et al. (2022)). We utilize multi-task learning to predict the final prediction label and one of the protected labels (e.g., gender). We save the model periodically in our experiments and implement Monte Carlo dropout, allowing us to get uncertainty estimations of these saved models.
Figure 1: Proposed Method Structure
3.3 Monte-Carlo Dropout
Modeling uncertainty with Monte Carlo dropout works by running multiple forward passes through the model with a different dropout mask every time. It draws weights from approximate posterior distributions. The posterior distribution is a way to summarize what we know about uncertain quantities in Bayesian analysis. It combines the prior distribution and the likelihood function, which tells you what information is in your observed data (the “new evidence”). In other words, the posterior distribution summarizes what we know after the data is observed.
Given a trained neural network model with dropout $f_{nn}$, to derive uncertainty for one sample $x$, the Monte Carlo algorithm collects the predictions of T inferences with different dropout masks. Here, $f_{nn}^{d_{i}}$ represents the model with dropout mask $d_{i}$. So a sample of the possible model outputs for sample $x$ is obtained as $f_{nn}^{d_{0}}(x),...,f_{nn}^{d_{T}}(x)$. By computing the average and variance of this sample, we obtain an ensemble prediction, which is the mean of the model’s posterior distribution for this sample and an estimate of the uncertainty of the model regarding $x$.
$predictive\>posterior\>mean:p=\frac{1}{T}\sum_{i=1}^{T}f_{nn}^{d_{i}}(x)$ (2)
[…]
More information on this can be found in the original publication on Monte Carlo dropout by Gal et. al Gal and Ghahramani (2016).
3.4 Non-Dominated Sorting
From the distribution of Monte Carlo models obtained, we saved their uncertainty scores and extracted the parameters of each model to utilize in a single (non-multi-task) prediction of our main target label. We trained each of these new models, collected their performance scores, and calculated the fairness scores (DIRs, difference in false negative scores, difference in false positive scores) from their produced predictions.
We implemented a non-dominated sorting algorithm for the performance scores and uncertainty scores, specifically the DIR scores of the model ensemble, to obtain the Pareto optimal set of models regarded as the "Pareto Front", that give the ideal trade-offs of performance and fairness. With the Pareto Front (or frontier), there are no better performing models in terms of performance and fairness scores, and there are only models where an improvement in one score leads to a reduction in the other. Non-dominated sorting allows one to search for optimal solutions for an optimization problem with multiple objectives. The algorithm in Figure 2 describes this process of non-dominated sorting. This algorithm takes two equally-sized lists, one is a list of performance scores for the models produced by the implementation of Monte Carlo dropout represented by X, and the second is fairness scores, specifically the DIR scores of said models, represented by Y. This method returns the elements which lie on the Pareto frontier, sorted in order. In our case, the objective is to maximize X, which is the performance score, and minimize the absolute difference between Y (DIRs), and 1.
Figure 2: Algorithm to Calculate Pareto Front
4 Data
We evaluated the effectiveness of our proposed method using three datasets, ADULT (individual income and demographic information) Kohavi et al. (1996), MIMIC-III (data from patients in critical care units) Johnson et al. (2016), and SNAPSHOT datasets (multi-modal physiological, behavioral, and survey data collected from 350 college students) Sano et al. (2015). See the Appendix A for more the details on the datasets.
4.1 ADULT Dataset
This dataset contains 14 categorical and continuous integer attributes from which we dropped columns with noisy data (over 100 missing values). We were left with 8 features: Age, Workclass, Education, Marital-status, Occupation, Relationship, Race, and Sex. Appendix A describes these features and their classes.
The prediction task for this dataset is a binary income/salary prediction, where a label of 1 represents participants who earn $>$ 50k and 0 for those who earn <=<=< = 50k in a year. We converted all the string features to integers, and we used 45222 of 48842 instances after dropping a total of 3,620 samples with noisy data. We utilized age, sex (gender), and race as the protected labels in our experiments.
For fairness analysis, we binarized the 3 protected labels to 0 as unprivileged and 1 as privileged classes. For the age label, 12% of the female class earned more than 50k in salary, while for the male class, 32% earned more than 50k. 36% of participants above the age of 40 earned more than 50k, while 17% of those below the age of 40 earned more than 50k. As for the race label, less than 13% of the participants who identified as Black, Amer-Indian-Eskimo, and "Other" earned more than 50k, while about 30% of those who identified as White or Asian-Pacific -Islander earned more than 50k. We chose the class corresponding to higher income/salary as the privileged class for each protected label. These privileged classes are also the majority classes in the dataset. For the age label, since the difference in the number of participants between classes is larger than the difference in average salary earned, we chose the class with more participants (less than or equal to 40) as the privileged class (1), while the class with fewer participants (over 40) as the unprivileged class (0). For race, Black(0), Amer-Indian-Eskimo(4), and other(2) became 0, and White(3) and Asian-Pac-Islander(1) became 1, and for sex, male became 1 and female became 0 as seen in table 1. We excluded the protected labels as features to feed into our prediction model. We split the data into 75% training and 25% testing sets to fit into the model.
Table 1: ADULT: Binarization of Protected labels
Label Privileged Class (1) Unprivileged Class (0)
Age Less than or equal to 40 years Greater than 40 years
Race White, Asian-Pac-Islander Black, Amer-Indian-Eskimo, other
Sex Male Female
4.2 MIMIC-III Dataset
With this dataset, we predicted in-hospital mortality (IHM) as a binary label. We determined IHM by comparing the patient’s date of death with hospital admission and discharge times. The resulting mortality rate is 13.23% (2,797 of 21,139 ICU stays). We processed data by reproducing the method in Harutyunyan et al. (2019). A full description can be found in Appendix A and the original publication Harutyunyan et al. (2019). To process the data for fairness evaluation, we binary-encoded each protected label and assigned them as privileged and unprivileged classes to analyze them using our chosen fairness metrics. As shown in Table 2, we encoded all participants with age $<=60$ as the unprivileged class (0), and those $>60$ as the privileged class (1). The reasoning behind this encoding is that the Centers for Disease Control and Prevention (CDC) defines an “older adult” as someone who is at least 60 years old for Disease Control et al. (2012). For the Gender label, male was encoded as the privileged class (1) due to its higher number of samples and because female patients appear to have a slightly higher IHM risk, making the female class the unprivileged class (0). We classified the participants according to whether they have private or government-assisted insurance. Those with private or self-care insurance represented the privileged class (1), and those with Medicare, Medicaid, and Government represented the unprivileged class (0). The reasoning behind this is the connection between government-aided insurance and socioeconomic conditions and age as mentioned earlier in this section. For marital status, we categorized married patients, those with life partners as privileged, and all others (no partner) as unprivileged. We did this in accordance with research showing that unmarried patients, including those who are divorced, separated, widowed, or never married, show elevated rates of certain health complications and death compared to those with partners Dhindsa et al. (2020); Johnson et al. (2000).
Table 2: MIMIC-III: Binarization of Protected labels
Label Privileged Class (1) Unprivileged Class (0)
Age Greater than 60 years Less than or equal to 60 years
Gender Male Female
Insurance Private, other Medicare, Medicaid
Marital Status Married, Life partner Single, Widowed, Divorced, Separated
4.3 SNAPSHOT Dataset
From the SNAPSHOT data, we predicted 4 self-reported mood labels, namely morning and evening happiness and calmness levels, using physiological features, mobile phone usage, and weather. We utilized the previous seven (7) days of data to predict the current day’s label. These labels appear as a score of 1 to 100 in the data set, and we binarized them with all scores less than or equal to 50 to mean sad or stressed and above 50 means happy or calm.
[…]
Some participants do not have race and ethnicity information, and we dropped samples for such participants when running experiments for race and ethnicity labels. In the end, we dropped 314 samples from the training data, reducing it from 4030 to 3716 samples, and the test samples were reduced from 1866 to 1715. We split the data into 70% training and 30% testing sets for all experiments to fit into the model.
Figure 3: SNAPSHOT Demographic Distribution
5 Experiments
5.1 Baseline Models and Fairness Analyses
The first part of our experiments involved developing baseline models that make the target predictions for all datasets (income for ADULT, IHM for MIMIC-III, and 4 happiness and calmness labels for SNAPSHOT). Next, we analyzed our prediction results for fairness by calculating the DIR and equalized odds for predictions against the protected labels for all datasets, respectively. We conducted these analyses to determine whether and where the models were introducing bias and ensure that we only apply our bias mitigation method to aspects of the data that are actually biased.
For the ADULT dataset, we built a Convolutional Neural Network (CNN) with 2 dense layers and a dropout later with a dropout rate of 0.25 to predict binary income labels; $<=50$k as 0 and $>50$k as 1. We used a 3-fold cross-validation method and binary cross entropy as the loss function with an Adam optimizer. The structure of this model is shown in Appendix B
For the MIMIC-III dataset, we used a Long-Short Term Memory (LSTM)-based model proposed in Harutyunyan et al. (2019). LSTM is a Recurrent Neural Network (RNN) designed to capture long-term dependencies in sequential data Harutyunyan et al. (2019). An RNN is a type of Artificial Neural Network (ANN) with recurrent connections capable of modeling sequential data for sequence recognition and prediction Bengio et al. (1994). To implement our model, we resampled the time series data into regularly spaced intervals. In the case of multiple measurements of the same variable in the same interval, we utilized the value of the last measurement. We used a 48-hour window for prediction, enabling the detection of patterns that may indicate changes in patient acuity, as proposed by Harutyunyan et al. (2019). We selected the channel-wise LSTM model without deep supervision for our study based on its superior reported area under the receiver operating characteristic (AUROC) performance among the five developed non-multitask models by Harutyunyan et al. Harutyunyan et al. (2019). AUROC is a performance metric that demonstrates a model’s ability to discriminate between cases (positive examples) and non-cases (negative examples.) Hanley and McNeil (1982). We also utilized the area under Precision-Recall (PR) curve (AUPRC), a single number that summarizes the information in the PR curve. AUPRC is a useful performance metric for imbalanced data in a problem setting where finding positive examples is essential. Particularly in medical datasets, the AUPRC is frequently smaller in absolute value than the AUROC since the fraction of positives is often less than 0.5, meaning that AUPRC has a lower baseline value than AUROC Ozenne et al. (2015).
This selected model is a modified version of the standard LSTM in which all variables are first independently pre-processed with an individual bidirectional LSTM layer instead of working directly on the entire matrix of clinical events as usual. Bidirectional LSTM (BiLSTM) is a recurrent neural network where, unlike standard LSTM, the input flows in both directions and can utilize information from both sides Graves and Schmidhuber (2005). Our network consists of a bi-directional layer, an LSTM layer with 16 neurons, and a dropout layer with a dropout rate of 0.3 (structure shown in Appendix B). We used a batch size of 8, binary cross entropy as the loss function and Adam optimizer. We executed this for 100 epochs, reproducing the experiments from Harutyunyan et al. (2019).
The baseline model we used for predicting the labels for the SNAPSHOT dataset is a GRU model with 2 GRU layers and 2 dropout layers with a dropout rate of 0.5, 30 neurons and focal loss with a gamma value of 2, and an alpha value of 4. This model predicts 4 labels, namely: "morning happiness", "morning calmness", "evening happiness", and "evening calmness". We trained this model for 300 epochs with a batch size of 1000 and a focal loss function since there is a significant class imbalance in the target label for this dataset. Focal loss is an extension of the cross-entropy loss function that would down-weight easy examples and focus training on hard negatives Lin et al. (2017). So it reduces the loss contribution from easy examples and increases the importance of correcting misclassified examples. The structure of this model is also shown in Appendix B. With 8 protected labels, which are gender, race, ethnicity, openness, conscientiousness, extraversion, agreeableness, and neuroticism, and 4 target labels, we conducted 32 sets of experiments to analyze the bias in this dataset.
After performing fairness analysis on each of baseline model, we selected the protected labels by which our models showed the most bias and tested our methods on all datasets. For all fairness analyses in our experiments, we utilized a fairness toolkit developed by IBM - AI Fairness 360 Bellamy et al. (2019).
5.2 Proposed Method Implementation
After running the multi-task networks and obtaining an ensemble of models using the parameters described in Appendix B, we calculated the uncertainties of these models. We computed the uncertainties by calculating the sample variance of the different forward passes using Equation 3. We built single-task models using the saved model weights to predict the target labels: income for ADULT, IHM for MIMIC-III, and the 4 mood-related labels for SNAPSHOT. With the predictions obtained from these models, we calculated the DIRs. These steps provided us with distributions of 100 performance scores and 100 DIR scores. We implemented the non-dominated sorting algorithm introduced in 3.4 to obtain the sets of Pareto optimal solutions for the experiments for each dataset. This algorithm computed an output containing the models with the best trade-off between the 2 scores (performance and DIR), and for which there were no better-performing models. Any model can be chosen from these models according to what is being prioritized for the particular problem type and dataset. We plotted the performance scores against the DIR scores to get a visual representation of their relationship.
We also conducted further analysis to determine whether or not there is a relationship between uncertainty and fairness. For this, we explored the relationship between the uncertainty scores and fairness scores (DI ratio) of the model ensemble produced by the Monte Carlo method.
5.3 Reweighing Method Implementation
We compared our proposed method to a standard bias mitigation method called Reweighing. Reweighing is a pre-processing bias mitigation technique. It involves applying appropriate weights to different tuples in the training dataset to make it discrimination free with respect to the protected label Park et al. (2021). It generates weights for the training examples in each (group, label) combination differently to ensure fairness before classification. Bellamy et al. (2018). See the details of Reweighting in Appendix B.
We applied the AI Fairness 360 implementation of this method to our processed data for all datasets and passed the weight vectors to our baseline model for all 3 datasets and calculated their fairness metric scores to compare to our proposed method.
5.4 Saliency Maps
To better understand the influence of the input features on our final predictions and also the difference our proposed method makes on how the networks determine what features to prioritize for prediction, we utilized a saliency map technique Simonyan et al. (2013) to visualize the importance of model weights on our input time axis and features. According to the predicted class $c$, for example, the decision-making process of the model can be represented as $S_{c}(I)=w_{c}^{T}I+b_{c}$, where $I,S_{c}(I),w,b$ are the model input, output, weights, and bias, respectively. The purpose of this method is to calculate the model weights on the input layer by gradients, for example, $w=\frac{\partial S_{c}}{\partial I}|_{I_{0}}$. The calculated $w$ represents the model saliency associated with the input layer. In this study, we fetched the saliency maps for all the test samples and calculated the average saliency map to develop the general intuition of important features and time steps.
We calculated the feature importance for our experiments according to the formulas in this section and plotted them in the figures in the results section.
6 Results
6.1 ADULT Dataset
The initial analysis before running a model showed that all 3 protected labels biased the data. For example, for the age label, the DIR score was above 2 for both the training and testing data; for sex, it was approximately 0.36 for both, and for race, it was approximately 0.47 for both. After running the base model, this trend continued, where the model introduced even more bias by age, making the DIR above 3 and below 0.1 for sex, At the same time, for race, it stayed similar to the initial analysis with a score of approximately 0.5.
Since our analyses on the baseline model showed bias by all 3 chosen protected labels, we experimented with our proposed method on all 3 protected labels. After implementing the proposed method and running the non-dominated sorting algorithm on the models we obtained, the Pareto fronts the method produced can be seen in Figures 4. To give a better, more linear visual representation of our results, we calculated the absolute value of [$1-$DIR Score] for the models to plot against the accuracy. The lines on the plots represent the Pareto fronts which contain the models that performed best to choose from. For all of the experiments on this dataset, we opted for the models that provided us the best fairness each time due to the insignificant differences in performance (accuracy) between the models (they all fall between 0.77 and 0.8). Therefore, we chose the models enclosed in the green circles to represent our final models.
Figure 4: ADULT: Pareto Front Plot for Age, Race, and Sex Labels
We also utilized the reweighing technique to improve fairness in these models as a comparison to our proposed method. We compared our final performance results of the base model, reweighing, and proposed method in Figure LABEL:adult_di_acc_plot. The reweighing technique slightly improved the fairness scores, with minimal impact on the performance scores of the models. For the protected age label, reweighing improved the DIR score by 0.84, for sex by 0.159, and for race by 0.0067 which we deemed not significant. The chosen model from our proposed method performed better than reweighing in terms of fairness, with an improvement in the DIR score of the age label by 1.28, sex by 0.32, and 0.042 for the race label. However, this improvement in fairness came at a slight performance cost, where the proposed method caused an average decrease of 1.7% in accuracy for all 3 labels. The reweighing method had an even lower performance cost, improving the accuracy score for the sex label. These results are visually represented in Figure 5 where the markers in blue represent the baseline models, red represents models implemented with reweighing technique, and green represents the models produced by our proposed method. In this figure, the x-axis represents an absolute value of the DIR score subtracted from 1 for an easier representation of our results.
Figure 5: ADULT: DIR and Accuracy Scores of Models
After calculating the feature importance for the baseline model and the models produced by our experiments for age, race, and sex labels, the scores we obtained showed us that our method redistributed weights to features less relevant to the protected label. This change is demonstrated in Figure 6. These figures show that the workclass feature holds the most weight across the board for all labels and remains equally important after the proposed method implementation. On the other hand, the relationship feature holds no weight in any of the prediction tasks across all experiments. The changes in importance are hence demonstrated more on the 3 remaining features: education, marital status, and occupation. Furthermore, the relative importance of these features changes across the experiments according to the demographic label we tested our proposed method on. For example, for the age label, the least important features are occupation, marital status, and education in that order, and the proposed method implemented with age shifted weight importance to these features. This trend is observed for the other labels of race and sex.
Figure 6: ADULT: Feature Importance Comparison Between Protected and Target Labels
6.2 MIMIC-III Dataset
The baseline model for IHM prediction using the MIMIC-III dataset resulted in Accuracy = 0.866, AUROC = 0.603, and AUPRC = 0.181.
Our analysis of the MIMIC-III data showed that the baseline model is only biased by marital status out of the 4 protected labels we analyzed, with a DIR score of 1.3. Regarding age, gender, and insurance labels, the baseline model showed unsubstantial bias where the DIRs calculated for each fell within the acceptable range of 0.8-1.2, as shown in Table 3. The differences in false negative scores are also significantly low across all the protected labels.
Table 3: MIMIC: Fairness Scores for Baseline Model
Label DI Ratio Score Diff in FN Scores
Age 1.008 -0.006
Marital status 1.308 0.005
[…]
Insurance 1.021 -0.004
Since our analyses showed bias by marital status, we experimented with our proposed method using our target label of IHM and marital status. After implementing the proposed method and running the non-dominated sorting algorithm on the models we obtained, the Pareto fronts the method produced can be seen in Figure 7 for DIR plotted against the chosen metric AUROC. To provide a better, more linear visual representation of our results, we calculated the absolute value of [$1-$DIR Score] for the models to plot against the accuracy as we did in 6.1. The lines on the plots represent the Pareto fronts which contain the models that performed best to choose from for each of the performance scores. Our proposed method improved the 3 performance scores for IHM prediction across the 4 protected labels. The general accuracy for the models on the Pareto front showed an average improvement of 1.74% from the baseline model. The AUROC score improved by an average of 0.20, and the AUPRC score by 0.197. In this experiment, we focused on prioritizing the AUROC score, as mentioned in section 5.1 because AUROC is prioritized for this dataset in previous studies Harutyunyan et al. (2019), and the accuracy variability for the models in the Pareto front is low (approximately 0.65%). We selected our final model from the AUROC vs. DI score Pareto plot (Figure 7). From this plot, the 2 circled models represent the optimal solutions from our method. The performance scores for these 2 models are represented in Table 4. We chose model 40 as our final solution since it provided an almost equal improvement in DIR as model 23, with a more significant improvement in AUROC, AUPRC, and accuracy.
Implementing the reweighing bias mitigation technique showed similar prediction performance when compared to our proposed method (as seen in Table 4), improving the AUROC, AUPRC, and accuracy slightly less than the proposed method when compared to the baseline. Reweighing improved the DIR score of the baseline model by 0.12 but did not perform as well as our proposed method. Regarding the difference in false negative scores, the reweighing technique performed worse than the baseline.
Table 4: MIMIC: Performance and Fairness of Models from Proposed Method, Reweighing Method, and Baseline Model
Score Model 23 Model 40 Reweighing Baseline Model
Accuracy 0.879 0.883 0.883 0.866
AUROC 0.771 0.797 0.777 0.603
AUPRC 0.354 0.367 0.332 0.181
DI Ratio 0.99 1.043 1.188 1.308
Diff in FN 0.0199 0.005 -0.05 0.005
Figure 7: MIMIC-III: Pareto Front Plot for AUROC vs DI Ratio Score
6.3 SNAPSHOT Dataset
After performing the initial set of 32 fairness analyses on the SNAPSHOT data, our results showed that the data itself did not contain substantial bias by any of the 8 protected labels we tested by. The scatter plot in Figure 8 shows that all scores fall within the acceptable range of 1.2 and 0.8 for the DIR score, except for the experiment with Race as a protected label and evening-sad-happy as the target label (specified by the red circle on the figure). For the 4 target labels, morning-sad-happy, morning-stressed-calm, evening-sad-happy, and evening-stressed calm, the F1 performance scores fell from approximately 0.7 to 0.8.
Figure 8: SNAPSHOT: Scatter Plot of DIR Scores from Baseline Model
We ran our proposed method on Race as a protected label and evening-sad-happy as the target label and obtained accuracy scores in the range of 0.76 to 0.77 and DIR scores between 0.80 and 0.88 for the model ensemble produced. We chose the model with the best DIR score of 0.882 as shown in Figure 9. This is an improvement of approximately 0.1 in fairness from the baseline model with a DIR score of 0.78. In terms of performance, this method did not affect the F1 score when compared to the baseline model. Both have F1 scores of approximately 0.76. Using the reweighing technique, we obtained a perfect DIR score of 1, but that came at a high cost of performance, with the F1 score dropping to 0.39, a loss of 37% in performance.
Figure 9: SNAPSHOT: Pareto Front Plot for Race vs Evening-sad-happy Label
6.3.1 Data Manipulation on SNAPSHOT
After analyzing the SNAPSHOT baseline model for bias and our results showed that the data was only biased by the Race label for evening-sad-happy label prediction; we performed some data manipulation to create a more significant imbalance for the gender label in order to determine if this would decrease the fairness performance and to test our proposed method on it. Originally, the data contained samples from 181 male and 132 female participants. There was a total of 4030 train and 1866 test samples, and we removed the samples from 75 participants that identified as female from the training data, making it a total of 2660 samples for training.We removed participants with 14 or more samples in the train data (experiment S-1). We also experimented with removing samples from participants with 5 or more samples in the train data (experiment S-2) and those with 30 or more samples (experiment S-3), leaving 2349 and 3685 samples, respectively. The motivation is to induce an artificial data bias, otherwise known as a negative legacy, into our model by creating a higher imbalance in the representation of male and female participant samples in our training data. With this manipulation, we expect our model to not generalize well on the test data that still contains a comparable number of female samples as hypothesized by Gorrostieta et al. (2019), thereby inducing some bias in its predictions.
After running the base model for all 4 target labels and analyzing fairness for all 3 experiment settings, we obtained similar performance scores. The fairness scores for experiments S-1 and S-2 were similar, with only the DIR score of the target label of evening-sad-happy occurring outside of the acceptable bounds of 0.8-1.2. In contrast, the fairness scores for experiment S-3 did not change significantly from the original SNAPSHOT experiment with no data manipulation. The F1 and fairness metric scores for experiment S-2 are listed in Table 5. As seen in the table, the DIR score of the target label of evening-sad-happy (the gray highlighted cell) fell outside the acceptable bounds of 0.8 and 1.2. All other fairness scores fell within these acceptable bounds.
Table 5: SNAPSHOT: F1 and Fairness Scores for Altered Data Experiments
Label F1 Score DI Ratio Score Diff in FN Diff in FP
Morning-Sad-Happy 0.749 1.153 -0.041 0.0979
Morning-Stressed-Calm 0.803 1.045 -0.009 0.073
Evening-Sad-Happy 0.798 1.291 -0.123 0.088
Evening-Stressed-Calm 0.788 1.0536 -0.0281 0.048
After running our proposed bias mitigation algorithm on the label combination of evening-sad-happy and gender on the manipulated data, we obtained an ensemble of models whose performance and DIR scores are plotted in Figure 10. On average, the models had an F1 score of 0.74 which is a 5.8% drop in performance from the baseline evening-sad-happy model. The average DIR score of the 9 models on the Pareto front is 1.034, which is well within the acceptable bounds of 0.8-1.2 and an overall improvement of 0.257 from the base DI ratio score of 1.291. We chose models 89 and 43 (green and red circles in Figure 10) as potential final models. Model 89 has the best DI ratio score of 1.001, which is almost the ideal score of 1, but it comes with a diminished F1 score of 0.722, a 7.6% drop in performance. It differs in False Positive and False Negative rates scores of -0.082 and -0.084, respectively. Model 43 has a DI ratio score of 0.995, with slightly better performance (0.733) than model 89.
Figure 10: SNAPSHOT: Pareto Front Plot for Gender vs Evening-sad-happy Label with Altered Data
7 Discussion
7.1 Fairness and Performance
With the current re-occurring limitations that bias and fairness research in machine learning face, our study aims to develop an effective, easily transferable bias mitigation technique, with an option of determining the tradeoff regarding fairness and performance of the model. Our proposed method improved fairness when tested on models developed for three datasets, ADULT, MIMIC-III, and SNAPSHOT, with minimal depreciation in performance and improvement in the case of MIMIC-III. Compared to the standard bias mitigation technique of reweighing, our proposed method performed better in significantly improving fairness for the ADULT and MIMIC-III experiments and slightly improved fairness for the SNAPSHOT dataset, but at no performance cost, while reweighing diminishing the F1 score by 37%.
From our experiments in predicting income with the ADULT dataset, our analyses showed us that the baseline model prediction of income was biased based on age, sex, and race, which corresponds with analyses from prior works (Le Quy et al. (2022); Lahoti et al. (2020); Kearns et al. (2019); Friedler et al. (2019); Zafar et al. (2017)). Implementing our proposed method provided a distribution of models that improved the DIR scores to within the ideal threshold of 0.8 and 1.2. This method performed better than the reweighing technique in terms of fairness, leading to an average of 1.7% decrease in accuracy for all three protected labels.
The LSTM model for IHM prediction using the MIMIC-III dataset was biased by marital status, not by age, gender, or insurance type. These results partially correlated with the fairness analysis performed in Röösli et al. (2022), where they found that gender was unaffected in terms of model fairness through their classification parity assessment Röösli et al. (2022). In their work, they did not consider marital status as a possible measure of fairness. In this work, we include marital status as a criterion due to various research showing evidence of the possible influence of marital status on one’s health. Dhindsa et al. (2020) showed that unmarried patients, including those who are divorced, separated, widowed, or never married have an increased rate of adverse cardiovascular events when compared to their married counterparts, and Johnson et al. (2000) found that non-married individuals show elevated rates of death compared to married persons, from the National Longitudinal Mortality Study (NLMS).
After finding that the model was biased by marital status, we implemented our bias-reducing technique, which improved the fairness to an almost-optimal level with a DIR score of 1.04 and a significantly low difference in False Negative scores between the privileged and unprivileged groups. In the process of doing this, our method also improved performance scores, raising the accuracy by 1.7%, AUROC score by 19.4%, and AUPRC by 18.6%. Compared to the reweighing technique, our method improved both fairness and prediction performances. We attribute this improvement in performance to a reason mentioned in prior research. As shown by the test for fairness-utility tradeoff conducted in Deho et al. (2022), ensuring fairness does not always lead to a drop in utility. Instead, it could lead to enhanced performance under specific circumstances.
For SNAPSHOT, we analyzed baseline models that predicted 4 self-reported mood-related labels to determine if they were biased by 3 demographic labels, namely gender, race, ethnicity, and 5 personality-based labels, namely openness, conscientiousness, extraversion, agreeableness, and neuroticism. These analyses determined that the DIR scores from all our experiments were within the chosen acceptable range of 0.8 to 1.2, except for the model predicting evening-happy-sad when analyzed by race. Implementing our proposed method improved the DIR score to 0.88, an improvement of 0.1 from that of the baseline method. In terms of performance, our proposed method had little effect in this case. Implementing the reweighing technique provided a perfect DIR score of 1 but a considerable performance cost of 37%.
As our initial analyses on the SNAPSHOT dataset baseline models did not show substantial bias by the 8 protected labels we used, we manipulated data by removing samples from 75 participants identified as female from the training data to determine the performance of our proposed method further. After running the baseline model on this altered data and analyzing, it showed bias by gender for the evening-sad-happy label with a DIR score of 1.29, slightly above our chosen acceptable range. Running our proposed technique resulted in an average, almost ideal DIR score of 1.034, an average 5.8% drop in performance.
7.2 Improving Model Explainability
With this work, we propose a method that improves fairness in machine learning models and allows practitioners in various fields to interpret better the decision-making process of the models they utilize. We understand that reducing model bias may not be sufficient in eliminating fairness concerns in decision contexts associated with certain fields such as predictive healthcare, which can be complex, unintuitive, and often hard to explain Yang (2022). In cases like these, solutions must strive for better explainability to improve understanding of the model, increase the utility of its output, and produce better outcomes Yang (2022). This is the reason our work introduces a possible way to strive for more transparency in the process by demonstrating which features affect model prediction the most and how our proposed method influences the model to avoid disparate treatment of unprivileged groups. This also corresponds with the fact that variables or features can and do correlate with protected labels such as race, age, etc. As mentioned in Paulus and Kent (2020), the use of “high dimensional” or “black box” prediction techniques that are typically favored in the machine learning community is generally problematic given that these approaches can predict the protected label through other variables, whether or not the actual protected label is explicitly encoded. This concern is eliminated with our work since our method specifically works by relying less on features connected to the protected label.
In our experiments involving the ADULT dataset, our visualization in Figure 6 showed how our method shifted importance from features that highly correlate with a particular protected label to reduce its effect on the final target label prediction.
For the MIMIC-III experiments, the saliency maps for the baseline model and the model after our proposed method was implemented are also listed in Appendix C. The deeper red color in these figures represents features that hold more weight and, therefore, are more critical for target prediction. Evaluating the map for the baseline model gives us an insight into some of the features that carried the most weight for IHM prediction. For example, the Glascow coma scale (GCS) motor response for Abnormal flexion and extension appears to be highly correlated with IHM prediction. GCS is a tool used to assess a patient’s level of consciousness after a brain injury Jain and Iverson (2018). The GCS motor response score evaluates the patient’s ability to obey commands or move in response to a stimulus. Numerous studies have examined the relationship between the GCS motor response score and in-hospital mortality Yumoto et al. (2019); Sadaka et al. (2012); Settervall et al. (2011). Generally, lower motor response scores are associated with a higher risk of mortality which explains the relevance of these features in our IHM prediction model. Two other features that are shown to be highly correlated with the baseline IHM prediction are Fraction inspired oxygen (FiO2), which is a measure of the concentration of oxygen that a patient is receiving, usually through supplemental oxygen therapy Fuentes and Chowdhury (2020), and pH, which is a measure of the acidity or alkalinity of a patient’s blood Arias-Oliveras (2016). Evidence suggests that higher FiO2 levels may be associated with an increased risk of in-hospital mortality de Jonge et al. (2008), where higher FiO2 levels were associated with an increased risk of mortality in both adult and pediatric patients. There is also some evidence from prior studies that indicate an association between abnormal pH levels and an increased risk of in-hospital mortality Park et al. (2020a), where both low and high pH levels were associated with increased mortality in critically ill patients.
Some features the baseline saliency map shows to carry the least weight in the baseline IHM prediction include diastolic blood pressure, oxygen saturation, and temperature. After implementing our proposed method, the resulting saliency map shows weight shifts and better weight distribution among the features. However, oxygen saturation still carries little weight in the final IHM prediction.
For experiments on SHAPSHOT data, the saliency maps listed in Appendix C give us insights into the features most important for the initial evening-sad-happy target label prediction. The deeper red color in these figures represents features that hold more weight and are, therefore, more important for target prediction across the 7 timesteps. From the saliency maps for the baseline method, a few features show the highest weights in the prediction task, at the top of this list of features being no_sleep_24, phys_3H-10H-medStepsWeightedPeaks, SMS_0H-3H_total_num. Comparing these maps to those from the model produced by our proposed method, it can be noted that our proposed method distributed the weights more evenly among all the features in general, explicitly moving some of the importance away from the more highly-weighted features in the baseline model prediction.
This trend of shifts in feature weights and importance across our experiments agrees with our hypothesis that a shift in feature weight importance to features that correlate less with the protected labels reduces the possibility of the model amplifying or creating biased connections between the protected label and target label.
From our initial work in Zanna et al. (2022), we attributed the notable drop in performance in the previous experiments using our method to the fact that both the target label and chosen protected labels shared similar important features, meaning that shifting weights from these features would have a significant effect on the model’s ability to efficiently form a pattern for the target label, ultimately affecting its prediction. Our results in this work prove this hypothesis to be true. From the visualization of the feature importance seen for ADULT, MIMIC-III, and SNAPSHOT, the models could utilize features that did not highly correlate with the protected label and still maintain an acceptable level of performance for the target label.
The explainable characteristic of our work also eliminates the need for researchers to initially identify the undesired correlation between the target label and the protected label to be able to tackle bias in the model using general existing bias mitigation methods Wang et al. (2020). We hypothesize that our technique can be implemented as a preventative method without knowledge of prior bias in the model. However, this is only applicable to models where our method does not diminish the performance to a detrimental level.
7.3 Performance-Fairness Trade-off
With the non-dominated sorting part of our method, we can enhance the performance-fairness tradeoff using the concept of Pareto optimality from multi-objective optimization. With this, we are able to obtain a set of Pareto-optimal solutions that give us the ideal models to choose from. With this method, we are providing flexibility for the proprietor to choose what objective, whether fairness or performance, they would like to prioritize and to what extent for every given solution, which can vary according to application or type of data. In our work, we are able to utilize this feature of our method to obtain models with the highest fairness scores since the variability in performance scores was generally low across all experiments.
7.4 Limitations
One limitation of this research is the restriction to binary outcomes imposed by the tools and techniques developed so far for bias mitigation. The implementation of the existing fairness analysis algorithms provided in the AIF360 toolkit created by IBM requires a binary label dataset as the input, which in models with non-binary labels, might be a barrier to adequately analyzing fairness.
Another limitation is the general lack of an agreed-upon definition of fairness in the field Park et al. (2020b). In the area of Machine Learning, most attempts at proposing new fairness constraints for algorithms have come from the West, and most of the works use the same datasets and problems to show how their constraints perform, which means that there is still no explicit agreement on which constraints are the most appropriate for which problems Mehrabi et al. (2021). Even though fairness is an incredibly desirable quality in society, it can be surprisingly difficult to achieve in practice. With these challenges in mind, there needs to be more standardization regarding fairness definitions and a more rigorous evaluation of what definitions should be applied to different data and problem types. In the future, it will be interesting to determine how popularly-used definitions of fairness translate to models built for data from different domains and how effective they are for each of these models.
Working with deep learning models, we have also encountered limitations associated with the computational complexity of using Monte Carlo dropout. Using this method along with the aspect of saving and loading model weights results in a higher than normal computational cost and time.
8 Conclusion
Our proposed method introduces a technique for bias mitigation that utilizes feature weights. We demonstrate how this method applies to various kinds of data from 3 different jurisdictions, successfully improving the fairness in the models regardless of the complexity of the predictor’s model. Our method improves fairness in the ADULT, MIMIC-III, and SNAPSHOT datasets at a greater or comparable rate (in the case of SNAPSHOT) to Reweighing, which is a standard bias mitigation technique without sacrificing performance to a substantial level.
Our method also provides an avenue for better model explainability/transparency through saliency maps to elaborate the shifts in feature importance in model prediction tasks. This is a principal characteristic of our work, as transparency is necessary to gain the trust of professionals in various industries, such as health care to have them adopt the predictive models into their daily operations Yang (2022).
Appendix A
In this appendix we provide further description of the ADULT, MIMIC-III and SNAPSHOT datasets from Section 4:
A.1 ADULT Data Description
The ADULT dataset was drawn from the 1994 United States Census Bureau data and involves using personal details such as education level to predict whether an individual will earn more or less than \$50,000 per year Kohavi et al. (1996). This dataset has been extensively utilized in machine learning bias and fairness research Zafar et al. (2017); Friedler et al. (2019); Kearns et al. (2019); Du et al. (2021); Fabris et al. (2022). The extracted features from the dataset used in this work are summarized in table 6.
Table 6: ADULT: List of extracted features
Feature Classes
Age Continuous
Education
Preschool, 1st-4th, 5th-6th, 7th-8th, 9th, 10th, 11th, 12th,
HS-grad, Some-college, Assoc-voc, Assoc-acdm, Prof-school, Bachelors,
Masters, Doctorate
Marital Status
Married-civ-spouse, Divorced, Never-married, Separated, Widowed,
Married-spouse-absent, Married-AF-spouse
Occupation
Tech-support, Craft-repair, Other-service, Sales, Exec-managerial,
Prof-specialty, Handlers-cleaners, Machine-op-inspct, Adm-clerical,
Farming-fishing, Transport-moving, Priv-house-serv, Protective-serv,
Armed-Forces
Race White, Asian-Pac-Islander, Amer-Indian-Eskimo, Other, Black
Relationship Wife, Own-child, Husband, Not-in-family, Other-relative, Unmarried
Sex Female, Male
Workclass
Private, Self-emp-not-inc, Self-emp-inc, Federal-gov, State-gov,
Without-pay, Never-worked
A.2 MIMIC-III Data Description
The MIMIC-III (‘Medical Information Mart for Intensive Care’) data set is a large, freely-available database that contains de-identified health-related data associated with over 40,000 patients who stayed in critical care units of the Beth Israel Deaconess Medical Center between 2001 and 2012. MIMIC-III contains detailed information regarding the care of patients, which includes vital signs, medications, laboratory measurements, observations and notes charted by care providers, fluid balance, procedure codes, diagnostic codes, imaging reports, hospital length of stay, survival data, patient demographics and in-hospital mortality (IHM), laboratory test results, billing information, etc. The complete data description can be found in the publication Johnson et al. (2016).
The risk of mortality is often formulated as binary classification using observations recorded from a limited window of time following admission Harutyunyan et al. (2019), and for our experiments, we predicted IHM as a binary label, so the target label indicates whether the patient died before hospital discharge. We processed and cleaned the data according to the method in Harutyunyan et al. (2019), where there are two major steps in the data processing pipeline. First, the root cohort is extracted based on the following exclusion criteria: Hospital admissions with multiple ICU stays or ICU transfers are excluded to reduce any possible ambiguity between outcomes, and ICU stays. Patients younger than 18 are excluded as well. Lastly, event entries are only retained if they can be assigned to a hospital and ICU admission and are part of the list of 17 physiological variables used for modeling, namely: Capillary refill rate, Diastolic blood pressure, Fraction inspired oxygen, Glascow coma scale, eye opening, motor response, verbal response, glucose, heart rate, height, mean arterial pressure, Oxygen saturation, respiratory rate, Systolic blood pressure, temperature, weight, and pH Röösli et al. (2022). From the root cohort, the IHM cohort was extracted by further excluding all ICU stays for which "length of stay" is unknown or less than 48 hours or for which there are no observations in the first 48 hours. The complete process of patient filtering is described in Figure 11 Harutyunyan et al. (2019). This provided final training and test sets of 17,903 and 3,236 ICU stays, respectively. Similar to Harutyunyan et al. (2019), We determined IHM by comparing patient date of death with hospital admission and discharge times. The resulting mortality rate is 13.23% (2,797 of 21,139 ICU stays) Harutyunyan et al. (2019).
The MIMIC-III dataset has some missing data in some instances. For these, we imputed the missing values using the most recent measurement value if it exists and a pre-specified “normal” value otherwise. These pre-specified values are provided by the authors in Harutyunyan et al. (2019). For categorical values, we encoded them using a one-hot vector and standardized the numeric inputs by subtracting the mean and dividing by the standard deviation. We calculated the statistics per variable after the imputation of missing values. After inputting the missing values and standardizing, we obtained 17 pairs of time series for each ICU stay. We have a binary target label for IHM for each stay, determining whether or not a patient dies in the hospital.
Figure 11: MIMIC-III: Data Preparation Workflow
From this dataset, the demographic data we utilized in our experiments include age, gender, marital status, and insurance, referred to as the protected labels in this paper. Table 7 shows the distribution of patients, ICU stays, and IHM rates in our data based on the different demographic groups. The age distribution is skewed towards older patients, with over 80% above 50 years. As for ethnicity, two-thirds of the patients are non-Hispanic White. In terms of health insurance, According to Röösli et al. (2022), health insurance type is included as a socioeconomic proxy, and this led to us including it in our experiments. In general, three major insurance types are available in the United States: The public programs Medicare and Medicaid, and private insurance providers. Medicare is a federal social insurance program, and anyone over 65 years or with certain disabilities qualifies for it, whereas Medicaid provides health insurance to very low-income children and their families. Insurance data was therefore mapped to the four distinct categories Medicare, Medicaid, private and other insurance. In this dataset, roughly a third of the patients are covered by private insurance, and more than 50% are enrolled in Medicare. The overall IHM rate in the MIMIC-III dataset is 13.23%, and IHM increases with age. As for gender, female patients appear to have a slightly higher IHM risk, and there is also a large range of variability for ethnic and socioeconomic groups.
Table 7: MIMIC-III: Demographic Distributions
Patients n(%) IHM rate (%)
Totals 18,094 13.23
Age
0-17 0 (0.0) 0
18-29 782 (4.3) 5.6
30-49 2,680 (14.8) 9.3
50-69 6,636 (36.7) 11.1
70-89 7,043 (38.9) 16.5
90+ 953 (5.3) 21.8
Gender
Female 8,090 (44.7) 13.5
Male 10,004 (55.3) 13
Insurance
Medicare 10,337 (57.1) 15.3
Medicaid 1,489 (8.2) 10.3
Private 5,601 (31.0) 10.2
Other 667 (3.7) 11.6
Marital Status
Married 8,564 (45.3) 46.4
Single 4,422 (24.4) 19.3
Widowed 2,654 (14.7) 17.4
Divorced 1,119 (6.2) 5.4
Separated 194 (1.0) 1.1
Unknown (default) 302 (1.7) 1.1
Life partner 5 (0.03) 0
A.3 SNAPSHOT Data Description
The SNAPSHOT dataset includes multi-modal physiological, behavioral, and survey data collected from 350 college students in one university, totaling over 7500 days of data Taylor et al. (2017). The study was conducted between 2013 and 2017, within which 30 days of consecutive data were recorded for 228 participants and 90 days for 15 participants. The data recorded includes wrist-worn wearable sensor data (acceleration, skin conductance, and skin temperature), mobile phone data (call, SMS, screen on/off logs, GPS locations), weather data (obtained using DarkSky API Dark Sky (2015)), self-reported daily morning and evening well-being (non-numeric scales or mood, health, and stress later scored 0-100), genders, and Big Five Personality scores John et al. (1999).
From the SNAPSHOT data, we computed a total of 378 features, including 173 physiological features that include Electrodermal activity (EDA) features, skin conductance Level (SCL), skin temperature features, accelerometer features, 165 mobile phone features, 43 weather features from the raw data, and statistical features calculated from each of these classes of features. Tables 8, 9, 10, 11, 12, and 13 summarize these features. For more details on the features extracted, refer to Taylor et al. (2017).
Table 8: SNAPSHOT: Description of Sleep and Nap Related Survey Data
Features Description
State Score
State anxiety score
no_sleep_24
Number of times participant slept in 24 hours.
sleep_latency
Time taken to fall asleep
pre_sleep_activity
What activity the participant performed before going to sleep.
awakening
If the participant woke up through the night (yes or no)
awakening_occations
Number of awakenings
wake_reason
How the participant woke up (spontaneously, alarm, or disturbance).
count_awakening
How many times participant woke up through the night.
awakening_duration
How long was the participant awake for if they woke up through the night.
nap
If the participant took a nap (yes or no)
nap_occations
Number of naps
count_nap
How many times the participant took a nap.
nap_duration
How long the nap was.
first_event_none
If an event is scheduled that day (yes or no)
time_in_bed
How long the participant spent in bed for the day.
sleep_try_time_mins_since_midnight
What time the participant tries to go to sleep
wake_time_mins_since_midnight
Wake time in minutes since midnight
first_event_mins_since_midnight
first event time in minutes since midnight
presleep_media_interaction
If the participant has presleep media interaction (yes or no)
presleep_personal_interaction
If the participant has presleep personal interaction (yes or no)
positive_interaction
If the participant had any positive interactions with someone for the day (yes or no).
negative_interaction
If the participant had any negative interactions with someone for the day (yes or no).
Table 9: SNAPSHOT: Description of Activity Related Survey Data
Features Description
academic
If the participant attended any academic activities (yes or no).
count_academic
How many academic activities the participant attended.
academic_duration
How long the participant spent on academic activities
study_duration
How many hours participant studied for, outside of academic activities.
exercise
If the participants engaged in any exercise-based activities (yes or no)
exercise_occations
How many times the participant engaged in exercise-based activities.
exercise_duration
For how long the participant exercised.
extracurricular
If the participant attended any other extracurricular activities (yes or no).
count_extracurricular
How many extracurricular activities the participant attended in the day.
extracurricular_duration
How long the participant attended extracurricular activities for.
overslept
If the participant overslept and missed any scheduled events.
caffeine_count
Total servings of caffeine participant had for the day.
drugs
If the participant had any other drugs or medication besides (yes or no).
drugs_alcohol
If the participant had any alcohol (yes or no).
drugs_alert
If the participant had any drugs to keep them alert (yes or no).
drugs_sleepy
If the participant had any drugs that made them sleepy (yes or no).
drugs_tired
If the participant had any drugs that made them tired (yes or no).
Table 10: SNAPSHOT: Description of Electrodermal Activity and Skin Conductance Features Extracted from Wearable Sensor Data
Features Description
Electrodermal activity (EDA) Peak Features
Sum AUC
Sum of the AUC of all peaks for this period where the amplitude of the peak is calculated as the difference from base tonic signal
Sum AUC Full
Sum of AUC of peaks where amplitude is calculated as difference from 0
Median RiseTime
Median rise time for peaks (seconds)
Median Amplitude
Median amplitude of peaks ($\mu$S)
Count Peaks
Number of detected peaks
SD Peaks 30 min
Compute number of peaks per 30 minute epoch, take standard deviation of this signal
Med Peaks 30 min
Compute number of peaks per 30 minute epoch, take median of this signal
Percent Med Peak
Percentage of signal containing 1 minute epochs with greater than 5 peaks
Percent High Peak
Same as Percent Med Peak
Skin Conductance Level (SCL) Features
Percent Off
Percentage of period where sensor was off
Max Unnorm
Maximum level of un-normalized EDA signal
Med Unnorm
Median of un-normalized EDA signal
Mean Unnorm
Mean of un-normalized EDA signal
Median Norm
Median of z-score normalized EDA signal
SD Norm
Standard Deviation of z-score normalized EDA signal
Mean Deriv
Mean derivative of z-score normalized EDA signal ($\mu$S/second)
Table 11: SNAPSHOT: Description of Accelerometer and Skin Temperature Features Extracted from Wearable Sensor Data
Features Description
Accelerometer Features
Step Count
Number of steps detected
Mean Movement Step Time
Average number of samples (at 8Hz) between two steps (aggregated first to 1 minute, then we take the mean of only the parts of this signal occurring during movement)
Stillness Percent
Percentage of time the person spent nearly motionless
Sum Stillness weighted AUC
Sum the weights of the peak AUC signal by how still the user was every 5 minutes
Sum Steps Weighted AUC
Sum the weights of the peak AUC signal by the step count over every 5 minutes
Sum Stillness Weighted Peaks
Multiply the number of peaks every 5 minutes by the amount of stillness during that period
Max Stillness Weighted Peaks
Max value for the number of peaks multiplied by the stillness for any five minute period
Sum Steps Weighted Peaks
Divide number of peaks every five minutes by step count and sum
Med Steps Weighted Peaks
Average value of number of peaks divided by step count for every 5 mins
Skin Temperature (ST) Features
Max Raw Temp
Maximum of the raw temperature signal (°C)
Min Raw Temp
Minimum of the raw temperature signal (°C)
SD Raw Temp
Standard deviation of the raw temperature signal
Med Raw Temp
Standard deviation of the raw temperature signal
Sum Temp Weighted AUC
Sum of peak AUC divided by the average temp for every 5 minutes
Sum Temp Weighted Peaks
Number of peaks divided by the average temp for every 5 minutes
Max Temp Weighted Peaks
Maximum number of peaks in any 5 minute period divided by the average temperature
SD Stillness Temp
Standard deviation of the temperature recorded during periods when the person was still
Med Stillness Temp
Median of the temperature recorded during periods when the person was still
Table 12: SNAPSHOT: Description of Features Extracted from Weather API 1
Features Description
Sunrise
Time of sunrise (UTC)
Moon_phase
Moon phase value on a scale of 0-1(new moon-full moon)
Apparent_temp_max
Maximum apparent temperature of the day (°F)
Apparent_temp_min
Minimum apparent temperature of the day (°F)
Temperature_max
Maximum temperature of the day (°F)
Temperature_min
Minimum temperature of the day (°F)
Avg_cloud_cover
Percentage of sky covered by cloud on a scale of 0-1
Avg_dew_point
Average dew point temperature
Avg_humidity
Daily average value of humidity on a scale of 0-1
Avg_pressure
Average atmospheric pressure on the sea level (hPa)
Morning_pressure_change
Trinary value of pressure difference between midnight and noon (rising, falling, steady)
Evening_pressure_change
Trinary value of pressure difference between noon and midnight (rising, falling, steady)
Avg_visibility
Average visibility (meters)
weather_precip_probability
Precipitation probability
Temperature_rolling_mean
Rolling average of temperature
Temperature_rolling_std
Rolling standard deviation of temperature
apparentTemperature_rolling_mean
Rolling average of apparent temperature
apparentTemperature_rolling_std
Rolling standard deviation of apparent temperature
apparentTemperature_today_vs_avg_past
Difference between today’s apparent temperature and its rolling average
pressure_rolling_mean
Rolling average of pressure
pressure_rolling_std
Rolling standard deviation of pressure
apparentTemperature_today_vs_avg_past
Difference in today’s apparent temperature and its rolling average
pressure_rolling_mean
Rolling average of pressure
pressure_rolling_std
Rolling standard deviation of pressure
pressure_today_vs_avg_past
Difference between today’s pressure and its rolling average
cloudCover_rolling_mean
Rolling average of cloud cover
cloudCover_rolling_std
Rolling standard deviation of cloud cover
cloudCover_today_vs_avg_past
Difference between today’s cloud cover and its rolling average
humidity_rolling_mean
Rolling average humidity
humidity_rolling_std
Rolling standard deviation humidity
humidity_today_vs_avg_past
Difference between today’s humidity and its rolling average
Table 13: SNAPSHOT: Description of Features Extracted from Weather API 2
Features Description
windSpeed_rolling_mean
Rolling average of wind speed
windSpeed_rolling_std
Rolling standard deviation of wind speed
windSpeed_today_vs_avg_past
Difference between today’s wind speed and its rolling average
precipProbability_rolling_mean
Rolling average of precipitation probability
precipProbability_rolling_std
Rolling standard deviation of precipitation probability
precipProbability_today_vs_avg_past
Difference between current precipitation probability and its rolling average
sunlight
Duration of sunlight (sec)
quality_of_day
Quality of the day defined in terms of 8 categories in the range -4, 4: clear=4, partly-cloudy=3, cloudy=2, wind=1, fog=-1, rain=-2, sleet=-3, snow=-4
avg_quality_of_day
Average value for quality_of_day
precipType
Type of precipitation as integer: None=0, Rain=1, Hail=2, Sleet=3, Snow=4, Other=5
max_precip_intensity
Maximum Precipitation volume (mm)
median_wind_speed
Median wind speed of the day (meter/sec)
median_wind_bearing
Median wind bearing of the day (degrees)
Appendix B
In this appendix, we provide more information on the experimental process from section 5. We provide model structures and parameters, and elaborate on the Reweighting method.
B.1 Model Structures
Figure 12: ADULT: Income Prediction Model Structure
Figure 13: MIMIC-III: IHM prediction Model Structure
Figure 14: SNAPSHOT: Mood Labels Prediction Model Structure
B.2 Model Parameters/Implementation
We applied our bias-reducing multi-task learning method by developing MTL models to predict target and protected labels by which the baseline models are biased. We first trained the multi-task model for each of the combinations for all datasets. We assigned different loss weight ratios to the label pairs, prioritizing the prediction label over the protected label to ensure that the model prioritizes the prediction label and make certain that its uncertainty score is kept lower than the protected label prediction. For the ADULT dataset, the ratio was 4.5 to 0.25; for MIMIC-III, 5 to 0.5; and for SNAPSHOT, 6 to 0.25. We arrived at these combinations by experimenting with a few different combinations. For example, with the ADULT dataset, the uncertainty of the age label remained around 0.60 for varying ratios up to 4 to 1 and improved to 0.62 at ratios 4 to 0.25. For the SNAPSHOT dataset, we used Focal loss as the loss function for both predictions and assigned gamma values 2 and 0 for the chosen target and protected labels, respectively. We did this because setting gamma 0 reduces the relative loss for well-classified examples. We trained the MTL networks for 100 epochs for all the datasets and saved the weights at each epoch, making it a total of 100 weight combinations for each experiment run.
B.3 Reweighting
The Reweighting bias mitigation technique works by applying different weights to each group-label combination according to the conditional probability of label by the protected label.
[…]
Assuming the dataset D is unbiased, in the sense that T and Label are independent, the expected probability $P_{exp}(t\wedge+)$ would be:
$P_{exp}(t\wedge+):=t\times+$
where t is the fraction of objects having T=t and ‘+’ the fraction of tuples having Label == +. In reality, however, the actual probability
$P_{act}(t\wedge+):=t\wedge+$
might be different. If the expected probability is higher than the actual probability value, it shows the bias towards label ‘-’ for T=t. We assign weights to t with respect to label ‘+’. The weight will be:
$W(T=t|x(Label)=+):={P_{exp}(t\wedge+)\over P_{act}(t\wedge+)}$
This weight of t for label ’+’ will over-sample objects with T=t for the label ’+’. The weight of t for label ’-’ becomes:
$W(T=t|x(Label)=-):={P_{exp}(t\wedge-)\over P_{act}(t\wedge-)}$
When applied to for example the ADULT dataset taking sex as the protected label, to remove the dependency between sex (T) and the target label (Label), we calculate a weight for each data object according to its T and Label value. In this example where both T=sex and Label are both binary attributes, only 4 combinations between the values of T and Label are possible, i.e., T=female(f) or T=male(m) can have Label values ’+’ or ’-’. We can then calculate the weight of a data sample with T=f and Label ’+’. Let’s assume that approximately 50% samples have T=m and 60% have Label value ’-’. so the expected probability of the sample can be computed as:
$P_{exp}(Sex=m|x(label)=-)=0.50\times 0.60$
but its actual probability is 20%. So the weight W will be:
$W(Sex=m|x(Label)=-)={0.50\times 0.60\over 0.2}=1.5$
The weights for the other combinations are also calculated in a similar manner.
We assign to every tuple a weight according to its T and Label values. The balanced dataset is then created by sampling the original training data, replaced with the assigned weights. On this balanced dataset the dependency-free classifier is learned. The Reweighing technique can be seen as an instance of cost-sensitive learning Elkan (2001) in which, e.g., an object of label ‘+’ with T=t gets a higher weight, making an error for this object more "expensive". The pseudocode of the algorithm describing this Reweighing technique can be found in Calders et al. (2009) and Kamiran and Calders (2009).
Appendix C
In this appendix, we provide the saliency maps for the MIMIC-III and SNAPSHOT Experiments. Figures 15 to 23 correspond with the SNAPSHOT experiments. For these figures, the images on the left represent the saliency map for the initial baseline model for evening-sad-happy prediction, and those on the right are the maps for the model generated from our proposed method. Figure 24 was generated from the initial baseline model for IHM prediction for the MIMIC experiments, and Figure 25 for the model from our proposed method implementation with bias mitigation based on marital status.
Figure 15: SNAPSHOT: Comparison of Saliency Maps From Baseline Model and Model After Proposed Method Implementation for evening-sad-happy Label Prediction for 47 features-1.
Figure 16: SNAPSHOT: Comparison of Saliency Maps From Baseline Model and Model After Proposed Method Implementation for evening-sad-happy Label Prediction for 48 features-2.
Figure 17: SNAPSHOT: Comparison of Saliency Maps From Baseline Model and Model After Proposed Method Implementation for evening-sad-happy Label Prediction for 48 features-3.
Figure 18: SNAPSHOT: Comparison of Saliency Maps From Baseline Model and Model After Proposed Method Implementation for evening-sad-happy Label Prediction for 48 features-4.
Figure 19: SNAPSHOT: Comparison of Saliency Maps From Baseline Model and Model After Proposed Method Implementation for evening-sad-happy Label Prediction for 48 features-5.
Figure 20: SNAPSHOT: Comparison of Saliency Maps From Baseline Model and Model After Proposed Method Implementation for evening-sad-happy Label Prediction for 48 features-6.
Figure 21: SNAPSHOT: Comparison of Saliency Maps From Baseline Model and Model After Proposed Method Implementation for evening-sad-happy Label Prediction for 48 features-7.
Figure 22: SNAPSHOT: Comparison of Saliency Maps From Baseline Model and Model After Proposed Method Implementation for evening-sad-happy Label Prediction for 24 features-8.
Figure 23: SNAPSHOT: Comparison of Saliency Maps From Baseline Model and Model After Proposed Method Implementation for evening-sad-happy Label Prediction for 19 features-9.
Figure 24: MIMIC-III: Saliency Maps From Baseline Model for IHM prediction.
Figure 25: MIMIC-III: Saliency Maps From IHM Prediction Model after Proposed Method Implementation with Bias Mitigation by Marital Status.